2013 2014 2015 Assets 01/01           450,000              540,000 ? Liabilities 01/01           100

2013 2014 2015
Assets 01/01           450,000              540,000 ?
Liabilities 01/01           100,000              150,000           120,000
Equity 01/01           350,000              390,000 ?
CS 01/01           200,000              220,000           235,000
RE 01/01           150,000              170,000 ?
A 12/31           540,000              505,000           545,000
L 12/31           150,000 ?           115,000
E 12/31           390,000 ? ?
CS 12/31           220,000 ?           250,000
RE 12/31           170,000 ? ?
Net income             40,000 ? ?
Dividends             20,000                15,000             30,000

I am stuck I want to figure out the rest of them not just the answers but how to solve it.
Answer #1

Answers

  • All working forms part of the answer
  • Following equations are used to find missing figures:

>Asset = Liabilities + Equity

>Increase (decrease) in Retained Earnings = Net Income (Loss) – Dividends

>CS + RE = Total Equity.

  • Answers, with missing figures:

2013

2014

2015

1

Assets 01/01

$                           450,000.00

$             540,000.00

$                   505,000.00

2

Liabilities 01/01

$                           100,000.00

$             150,000.00

$                   120,000.00

3

Equity 01/01

$                           350,000.00

$             390,000.00

$                   385,000.00

4

CS 01/01

$                           200,000.00

$             220,000.00

$                   235,000.00

5

RE 01/01

$                           150,000.00

$             170,000.00

$                   150,000.00

6

A 12/31

$                           540,000.00

$             505,000.00

$                   545,000.00

7

L 12/31

$                           150,000.00

$             120,000.00

$                   115,000.00

8

E 12/31

$                           390,000.00

$             385,000.00

$                   430,000.00

9

CS 12/31

$                           220,000.00

$             235,000.00

$                   250,000.00

10

RE 12/31

$                           170,000.00

$             150,000.00

$                   180,000.00

11

Net income

$                             40,000.00

$               (5,000.00)

$                     60,000.00

12

Dividends

$                             20,000.00

$               15,000.00

$                     30,000.00

  • Working for above missing figures:

2013

2014

2015

1

Assets 01/01

450000

540000

505000

2

Liabilities 01/01

100000

150000

120000

3

Equity 01/01

350000

390000

=505000-120000

4

CS 01/01

200000

220000

235000

5

RE 01/01

150000

170000

150000

6

A 12/31

540000

505000

545000

7

L 12/31

150000

120000

115000

8

E 12/31

390000

=505000-120000

=545000-115000

9

CS 12/31

220000

235000

250000

10

RE 12/31

170000

=385000-235000

=430000-250000

11

Net income

40000

-5000 [Decrease in RE = $ 20,000 including decrease of $ 15,000 of dividend, hence Net Loss $ 5,000]

60000 [Net Increase in RE = $ 30,000 which include decrease by Dividends of $ 30,000, Hence Net Income = 30000 + 30000 = $ 60,000

12

Dividends

20000

15000

30000